You decide to put $10,000 in a money market fund that pays interest at the annual rate of 7.2%, compounding it monthly. You plan to take the money out after one year and pay the income tax on the interest earned. You are in the 25% tax bracket. Find the total amount available to you after taxes.
Answer #1

i = 7.2% / 12 = 0.6% per month

t = 12 month

Interest income = 10000 * (1+0.006)^12 - 10000

= 10000 * (1.006)^12 - 10000

= 10744.24 - 10000

= 744.24

Tax = 0.25 * 744.24 = 186.06

Total amount after taxes = 10000 + 744.24 - 186.06 = 10558.18